Project

General

Profile

NDOS Commissioning Meeting 2011/04/11

Attendance: Jon, Andrew, Kurt, Peter, Mark Messier, Mark Mathis, Sue, Jarek, Leon, Alena, Mat, Greg, Rick K., Athans

Agenda:

  1. NDOS update
  2. DAQ Update:
  3. DCS Update:
  4. DAQ update
  5. AOB

NDOS Status:

  • ~30 FEBs suddenly dropped out of the readout "recently". DSO scan from this morning indicates these APDs are very noisy. Related to recent high humidity?
  • How to track this in onmon/online?
    • DCM application can look for FEBs that have turned off.
    • Onmon should reset/make new plots for every subrun (Mark Messier will take a look at onmon while on shift tonight).
  • Plans for work on NDOS this week: Karen K will be working on filling this week.

DAQ:

  • dcm 3-1 was brought on last Friday and the process of including it into the data stream went pretty smoothly.
  • could not read out muon catcher, most likely because timing was not being sent through empty DCMs upstream. Could be an issue with the dcmapplication. Sue will look into this.
  • DAQ log messages: filling up disk. Needs better "monitoring".
  • Run Control did not roll over to the next run properly on Saturday. Jon will look into this.
  • DetectorPlots Package and First Ganglia-vs-DCS Plots (Peter S.)

DCS:

  • Andrew: Chad has installed a patch panel out at the detector. Also the computer in rack 1 was moved to rack 3 and is operational.
  • NDOS DAQ_DCS:the Test-Stand version (Athans)
    • Lots of work has gone into this, but there are several areas that need to be improved.
    • Athans is requesting that we schedule a time to try his code out on the detector; several people commented that this needs to be tested while at least one dcm is collecting data on the test-stand before we can consider running this on the detector.
    • Athans will be arriving at FNAL on Thursday morning. Jon strongly suggested that the DAQ Software and DCS groups meet sometime in the afternoon.

AOB

  • Since there is a collaboration meeting at the end of the week, this Thursday's commissioning meeting is canceled.